Grade 12Physical Chemistry

A 20ml sample of cuso4 was evaporated to dryness leaving 0.967g of residue . What was the molarity of the original solution (cu=63.5)

Komal kumari
8 Years agoGrade 12

1 Answer

Sandeep Bhamoo
ApprovedApproved Tutor Answer8 Years ago
Mass of CuSO4=0.967g. Vol. of soution=20ml=0.02l. Mole of CuSO4= 0.967÷159.5. Now molarity=mole÷volume =(0.967/159.5)÷0.02. =0.303N. Please do approve my answer if correct